OpenStudy (anonymous):

Differential Equation: y' = (1-2t)/y with initial Value of y(1) = -2 I get y = sqrt(2t - 2t^2 +4 ) but it says my answer is wrong. I've redone it a few times, and can't pick up where I did wrong.

myininaya (myininaya):

\[y'=\frac{1-2t}{y}\]

myininaya (myininaya):

looks like a separation of variables problem i'm excited! :)

myininaya (myininaya):

\[\frac{dy}{dt}=\frac{1-2t}{y}\]

myininaya (myininaya):

multiply y on both sides and multiply by dt on both sides to obtain \[y dy =(1-2t) dt\]

OpenStudy (anonymous):

Yup. Is seperation of variables :3

myininaya (myininaya):

now integrate both sides

OpenStudy (anonymous):

but I keep getting the same answer ;c

myininaya (myininaya):

\[\int\limits_{}^{} y dy= \int\limits_{}^{}(1-2t) dt\]

myininaya (myininaya):

\[\frac{y^2}{2}=t-t^2+C\]

myininaya (myininaya):

\[y^2=2t-2t^2+C\]

myininaya (myininaya):

2 times a constant is still a constant so instead of 2c i just wrote c

myininaya (myininaya):

now if t=1 then y=-2 so apply initial condition we obtain \[(-2)^2=2(1)-2(1)^2+C \] 4=2-2+C 4=C so we have \[y^2=2t-2t^2+4\]

myininaya (myininaya):

\[y=\pm \sqrt{2t-2t^2+4}\]

OpenStudy (anonymous):

Ah. plus minus. I forgot about that q-q Thanks :3

myininaya (myininaya):

np gj to you for doing everything correctly up into that point lol i mean you had one solution for y thats real good